# Fenglin Liu1, Jinge Wu211footnotemark: 1, Hongjian Zhou1, Xiao Gu1, Soheila Molaei1,&Anshul Thakur1, Lei Clifton3, Honghan Wu2,4, David A. Clifton1,5

### v1

@@ Acknowledgements
This work was supported in part by the Pandemic Sciences Institute at the University of Oxford; the National Institute for Health Research (NIHR) Oxford Biomedical Research Centre (BRC); an NIHR Research Professorship; a Royal Academy of Engineering Research Chair; the Well-come Trust funded VITAL project; the UK Research and Innovation (UKRI); the Engineering and Physical Sciences Research Council (EPSRC); the InnoHK Hong Kong Centre for Cerebro-cardiovascular Engineering (COCHE); and the Clarendon Fund and the Magdalen Graduate Scholarship.
@@
Author Contributions
F.L. conceived and designed the study. F.L. and J.W. contributed to the data analysis, implementations, and experiments, and prepared the manuscript. All authors contributed to the final manuscript preparation.
Competing Interests
The authors declare no competing interests.
Appendix A Method
We now introduce our method in detail. As shown in Figure 1 (a) and Figure 2, our RiskAgent introduces three LLM agents and one component, i.e., Environment, Decider, Executor, and Reviewer.
Environment
@@ Appendix A Method
Retrieval.
We first adopt the text-embedding-ada-002 model [48] to extract the embeddings $E_{T}$ of all tools $T$ using their tool metadata including both tool names and descriptions, with the dimension of 1,536, defined as follows:
@@
$S=\text{Cosine\_Similarity}(T,P)=\frac{T\cdot P}{\|T\|\|P\|}$ (3)
where $T\cdot P$ represents the dot product of the two vectors; and $\|T\|$ and $\|P\|$ are the Euclidean norms (i.e., $L_{2}$ norms) of the vectors. Therefore, we can set a threshold and retrieve all relevant tools with a similarity score greater than it.
Rank.
In our work, we propose ranking the tools using the calculated similarity scores $S$ and forwarding the top-$N$ tools with the highest similarity scores for the Decider to perform tool selection. We evaluate the performance of setting different values of $N$. Our preliminary results show that setting $N=5$ achieves a 99.8% recall score, meaning that in 99.8% of cases, the correct tool is included among the top-5 tools retrieved and ranked by our method. This proves the effectiveness of the embedding-based retrieval-ranking algorithm used.
@@
1. Output format MUST be: {"name": value}
2. Use EXACT "name" fields from schema as keys
@@ 4. Use ONLY values defined in schema options
ANALYSIS: "All stages processed correctly" / Stage_X: [ERROR] <error description>
@@ 2. If error found, report only the earliest error stage with justification
Instruction Fine-tuning
As we mentioned, enabling LLMs to accurately solve diverse medical tasks traditionally requires extensive medical data (ranging from 1 million to 80 billion tokens [15, 11, 49, 20, 21, 8]) and substantial computational resources that may be prohibitively expensive for fine-tuning.
Following the workflow established in Environments 1-5, we construct samples with explicit instructions and well-defined output formats. Each sample includes: (a) patient information, clinical queries, and candidate tools, with the expected output being the selection of the correct tool with justification; (b) patient information and tool schema, with the expected output being correctly formatted parameter extraction; (c) the original results computed by calculators with the sample question, with the expected output being a rephrased answer to the question; (d) outputs from step c along with questions and options, with the expected output being the final selected option; and (e) information from steps a-d, with the expected output being a comprehensive reflection (the instruction template is shown in Figure 6).
We take the LLaMA-3-8B LLM [35] as our backbone to train our model using parameter-efficient fine-tuning techniques (LoRA) [50], significantly reducing computational requirements while maintaining high performance across medical risk prediction tasks.
Hyper_parameter Value
@@
Experiment Settings
We perform model fine-tuning using LoRA[50] following the approach introduced using the LLaMA Factory framework [51]. The MedRisk benchmark dataset is divided in a 7:1:2 ratio for training, validation, and testing, respectively. All training is conducted using PyTorch’s DistributedDataParallel on 4$\times$×A100 (80GB) GPUs with BF16 precision. We use AdamW optimizer [52] with a cosine learning rate scheduler. The effective batch size is 32 (batch size 8 per device with 4 gradient accumulation steps). The detailed hyper-parameters for training are shown in Table 4.
During evaluation, for a fair comparison, we test all models on the same test set using a consistent protocol. We use a temperature of 0.6 and top-p of 0.9 for controlled randomness in generation. Each model receives identical input formatting, consisting of patient information, a clinical question, and four multiple-choice options. The evaluation metrics focus on accuracy, calculated as the exact match between model predictions and ground truth answers.
Ethical considerations
Our study was conducted on retrospective, de-identified data. For the public data, all necessary patient/participant consent has been obtained, and the appropriate institutional forms have been officially archived.

### v2

@@ Acknowledgements
DAC was funded by an NIHR Research Professorship; a Royal Academy of Engineering Research Chair; and the InnoHK Hong Kong Centre for Cerebro-cardiovascular Engineering (COCHE); and was supported by the National Institute for Health Research (NIHR) Oxford Biomedical Research Centre (BRC) and the Pandemic Sciences Institute at the University of Oxford. FL was funded by the Clarendon Fund and Magdalen Graduate Scholarship. HZ was funded by the Clarendon Fund, the Department of Engineering Science Studentship, and the Frederick Brodckhues Scholarship.
@@
Appendix A Framework
A.1 Environment
@@
Retrieval. We first adopt the text-embedding-ada-002 model (tex, 2022) to extract the embeddings $E_{T}$ of all tools $T$ using their tool metadata including both tool names and descriptions, with the dimension of 1,536, defined as follows:
@@
$S=$ $\text{Similarity}(T,P)$ (3)
$=$ $\text{Cosine\_Similarity}(E_{T},E_{P})$
$=$ $\frac{E_{T}\cdot E_{P}}{\|E_{T}\|\|E_{P}\|}$
where $(\cdot)$ represents the dot product of the two vectors; and $\|\cdot\|$ is the Euclidean norms (i.e., $L_{2}$ norms) of the vector. Therefore, we can set a threshold and retrieve all relevant tools with a similarity score greater than it.
Rank. In our work, we propose ranking the tools using the calculated similarity scores $S$ and forwarding the top-$N$ tools with the highest similarity scores for the Decider to perform tool selection. We evaluate the performance of setting different values of $N$. Our preliminary results show that setting $N=5$ achieves a 99.8% recall score, meaning that in 99.8% of cases, the correct tool is included among the top-5 tools retrieved and ranked by our method. This proves the effectiveness of the embedding-based retrieval-ranking algorithm used.
@@
1. Output format MUST be: {”name”: value}
2. Use EXACT ”name” fields from schema as keys
@@ 4. Use ONLY values defined in schema options
ANALYSIS: ”All stages processed correctly” / Stage_X: [ERROR] ¡error description¿
@@ A.2 LLM Agents
Appendix B Settings and Baselines
B.1 Settings
[…]
Hyper_parameters Value
@@ B.1 Settings
During evaluation, for a fair comparison, we test all models on the same test set using a consistent protocol. We use a temperature of 0.6 and top-p of 0.9 for controlled randomness in generation. Each model receives identical input formatting, consisting of patient information, a clinical question, and four multiple-choice options. The metrics focus on accuracy, calculated as the exact match between model predictions and ground truth answers.
B.2 Baselines
To comprehensively evaluate LLMs’ performance for risk prediction, we collect 19 different types of representative LLMs and agents. 1) Public LLMs: We collect multiple representative open-source public LLMs across different numbers of model parameters, i.e., Mistral-7B (Jiang et al., 2023), LLaMA-3-8B (AI@Meta, 2024), Mixtral-8x7B (Jiang et al., 2024), and LLaMA-3-70B (AI@Meta, 2024); We also collect three public medical LLMs: PMC-LLaMA-13B (Wu et al., 2024), Meditron-70B (Chen et al., 2023), and MedGemma (Sellergren et al., 2025). They are developed by fine-tuning general LLMs on a large corpus of medical datasets (Zhou et al., 2023) to learn rich medical knowledge, achieving improved performance in medical tasks. 2) Commercial LLMs: We further collect the currently most powerful commercial LLMs, GPT-3.5 (OpenAI, 2022), GPT-4 (OpenAI, 2023), GPT-4o-2024-08-06 (Hurst et al., 2024), o3-mini-2025-01-31 (gpt, 2025a), o1-2024-12-17 (Jaech et al., 2024), GPT-4.5 (gpt, 2025b), Claude-3.7-Sonnet (Anthropic, 2024a), Claude-4-Sonnet (Anthropic, 2024c), and Claude-4-Opus (Anthropic, 2024b). 3) AI Agents: Recently, AI agents (Wang et al., 2024; Qiao et al., 2024), which assign different LLMs to different roles to collaboratively perform tasks, have demonstrated better task performance than single LLMs in traditional domains such as gaming and programming. Therefore, for a comprehensive comparison, we further collect three well-known and recent AI agents, i.e., ReAct (Yao et al., 2023), BOLAA (Liu et al., 2023), and Chameleon (Lu et al., 2024). We adopt LLaMA-3-8B (AI@Meta, 2024) as the backbone of our system to re-implement them for a fair comparison.
Figure 6: Performance for RiskAgent-8B, GPT-4o, and o3-mini. In the boxplot, the central line indicates the median value, while the lower and upper boundaries represent the 25th (Q1) and 75th (Q3) percentiles, respectively. The whiskers extend up to 1.5 times the interquartile range (IQR). We can see that RiskAgent-8B not only consistently achieves the best results across all cases in different runs but also achieves lower perturbations (STD) than o3-mini in most cases, especially in risk prediction for cardiac.